OpenStudy (korosh23):

Determine all the co-terminal angles. Please wait for me until I draw all the required numbers. Thank you.

OpenStudy (korosh23):

\[\frac{ 11\pi }{ 6 }\] domain limit is -4pi smaller or equal to feda smaller than 4 pi \[-4\pi \le \Theta < 4\pi\]

OpenStudy (korosh23):

I have the answer I just need someone to check it for me.

OpenStudy (korosh23):

pi/6 7pi/6 13pi/6 19pi/6 23pi/6

OpenStudy (korosh23):

-23pi/6 *

OpenStudy (anonymous):

Well, that radian measure is the same as 345 degrees. Coterminal angles are angles who share the same initial side and terminal sides. An angle can have MULTIPLE coterminal angles. One question, are all of those angle choices negative?

OpenStudy (korosh23):

No, they are not a choice, that is my work. I did them, and I just need someone to check it for me.

OpenStudy (anonymous):

I apologize, the radian measure given is ACTUALLY 330 degrees. :D If your final answer is \[23\pi/6\], then you are CORRECT. Note that this cannot be negative. Simply add 360 degrees to ANY angle to get one of their coterminal angles. So, let's do that to 330: \[330+360= 690\] 690, converted to radian measure, is 23pi/6. :D

OpenStudy (anonymous):

Also note that another coterminal angles is: \[-\pi/6\]

OpenStudy (anonymous):

Sorry, ''angles'' is supposed to be ''angle''. :)

OpenStudy (korosh23):

Thank you. :)